OpenStudy (anonymous):

Solve 7|x-6|=49 ...Answer is x=13, correct?

OpenStudy (anonymous):

there is one more

OpenStudy (anonymous):

|x-6|=7 (x-6) can be 7 or -7 x-6=-7 x=-1

OpenStudy (anonymous):

Oh? Well i got that out of 3 other choices i know it cant be x=55 and x=-43

OpenStudy (anonymous):

Ohhhh kk i understand. Therei s a choice here answer D. x=13 or x=-1

OpenStudy (anonymous):

sorry my bad x-6=49/7=7 x=7+6=13 i am getting the answer as +13...

OpenStudy (anonymous):

No i get what your saying :)